OpenStudy (anonymous):

Find the y-intercept (b) using the formula y = mx + b and the points (2, 8) and (1, 3). Question 5 options: b = 2 b = -14 b = 0.6 b = -2 @iYuko

OpenStudy (anonymous):

these ones are hard for me lol..

iYuko (iyuko):

\[m=\frac{ y1-y2 }{ x1-x2 }\]

iYuko (iyuko):

8 is y1 3 is y2 2 is x1 1 is x2

iYuko (iyuko):

@kmc3490

iYuko (iyuko):

It will look like so: \[y-8=\frac{ 3-8 }{1-2 }(x-2)\]

iYuko (iyuko):

\[y-8=5(x-2)\]

iYuko (iyuko):

\[y-8=5x-10\]

iYuko (iyuko):

Final answer : \[y=5x-2\]

iYuko (iyuko):

The last option is the answer @kmc3490

OpenStudy (anonymous):

Okay thanks so much. I understand it now

iYuko (iyuko):

:)